In a slow year, Deutsche Burgers will produce 2.900 million hamburgers at a total cost of $4.300 million. In a good year, it can produce 4.900 million hamburgers at a total cost of $5.500 million a. What are the fixed costs of hamburger production? (Do not round intermediate calculations. Enter your answer in millions rounded to 1 decimal place.) Fixed cost million b. What is the variable cost per hamburger? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Variable cost per burger c. What is the average cost per burger when the firm produces 2 million hamburgers? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Average cost per burger d. What is the average cost per burger when the firm produces 3 million hamburgers? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Average cost per burger
Answer #1

a) Fixed Cost + 2,900,000*Variable Cost = 4,300,000
Fixed Cost + 4,900,000*Variable Cost = 5,500,000

Using both equation we get
2,000,000 * Variable Cost = 1,200,000
Variable Cost = 0.6 per hamburger

Fixed Cost = 4,300,000 -2,900,000*0.6 = 2,560,000

b) Variable Cost = 0.6 per hamburger

c) Total cost of 2,000,000 hamburgers = 2,560,000 + 0.6*2,000,000 = 3,760,000
Average Cost = 3,760,000/2,000,000 = 1.88

d) Total cost of 3,000,000 hamburgers = 2,560,000 + 0.6*3,000,000 = 4,360,000
Average Cost = 4,360,000/3,000,000 = 1.45

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well